OpenStudy (anonymous):

Dora can run 3/5 mi in 3/20 h. How far can she run in 1 h? mi

OpenStudy (anonymous):

so do we use the same form

OpenStudy (agl202):

\frac{\frac{3}{5} \ mi}{\frac{3}{20} \ h}=\frac{3}{5} \times \frac{20}{3} \ \frac{mi}{h}= \frac{20}{5} \ \frac{mi}{h}=4 \ \frac{mi}{h} = \frac{4 \ mi}{1 \ h} In 1 h she can run 4 miles.